P(-1,-4) lies on the curve y=x^2+5x.

If Q is the point (x,x^2+ 5x), find the slope of the secant line PQ or x= -0.5
1
Expert's answer
2020-06-29T18:53:25-0400

Consider the curve "y=x^2+5x" and a point "P(-1,-4)" lies on the curve.


For "x=-0.5" , the coordinates of point "Q(x,x^2+5x)" is,


"Q=(-0.5,(-0.5)^2+5(-0.5))=(-0.5,-2.25)"


The slope of the secant line joining the points "P" and "Q" is evaluated as,


Slope"(m)=\frac{y_2-y_1}{x_2-x_1}"


"=\frac{-2.25-(-4)}{-0.5-(-1)}"


"=\frac{-2.25+4}{-0.5+1}"


"=\frac{1.75}{0.5}"


"=\frac{7}{2}"


"=3.5"


Therefore, the slope of the secant line "PQ" is "m=3.5"